solve the equation for x by factoring the left-hand side.

x squared + 4x + 4 = 7

Answer by checkley71(8403)

X^2+4X+4=7
(X+2)^2=7
X+2=SQRT7
X+2=2.6575
X=2.6575-2
X=.6575 ANSWER.